### v1

# Breaking the Manual Annotation Bottleneck: Creating a Comprehensive Legal Case Criticality Dataset through Semi-Automated Labeling

Abstract
Predicting case criticality helps legal professionals in the court system manage large volumes of case law. This paper introduces the Criticality Prediction dataset, a new resource for evaluating the potential influence of Swiss Federal Supreme Court decisions on future jurisprudence. Unlike existing approaches that rely on resource-intensive manual annotations, we semi-automatically derive labels leading to a much larger dataset than otherwise possible. Our dataset features a two-tier labeling system: (1) the LD-Label, which identifies cases published as Leading Decisions (LD), and (2) the Citation-Label, which ranks cases by their citation frequency and recency. This allows for a more nuanced evaluation of case importance. We evaluate several multilingual models, including fine-tuned variants and large language models, and find that fine-tuned models consistently outperform zero-shot baselines, demonstrating the need for task-specific adaptation. Our contributions include the introduction of this task and the release of a multilingual dataset to the research community.
Breaking the Manual Annotation Bottleneck: Creating a Comprehensive Legal Case Criticality Dataset through Semi-Automated Labeling
Ronja Stern 1thanks:    Equal contribution.                        Ken Kawamura 5
Matthias Stürmer 1,2                        Ilias Chalkidis 4                        Joel Niklaus 1,2,3∗
1University of Bern 2Bern University of Applied Sciences 3Stanford University 4University of Copenhagen 5Independent Scholar
FSCS
Federal Supreme Court of Switzerland
SCI
Supreme Court of India
ECHR
European Convention of Human Rights
ECtHR
European Court of Human Rights
SCOTUS
Supreme Court of the United States
SPC
Supreme People’s Court of China
SJP
Swiss-Judgment-Prediction
ASO
Almost Stochastic Order
ILDC
Indian Legal Documents Corpus
US
United States
EU
European Union
NLP
Natural Language Processing
ML
Machine Learning
LJP
Legal Judgment Prediction
SJP
Swiss-Judgment-Prediction
PJP
Plea Judgment Prediction
BERT
Bidirectional Encoder Representations from Transformers
LSTM
Long Short-Term Memory
GRU
Gated Recurrent Unit
BiLSTM
Bidirectional Long Short-Term Memory
CNN
Convolutional Neural Networks
PLM
pre-trained Language Model
LM
Language Model
RTD
Replaced Token Detection
CLT
Cross-Lingual Transfer
HRL
high resource language
LRL
low resource language
POS
Part-of-Speech
SLTC
Single Label Text Classification
MLTC
Multi Label Text Classification
TC
Text Classification
NLU
Natural Language Understanding
IR
Information Retrieval
NER
Named Entity Recognition
NLU
Natural Language Understanding
QA
Question Answering
NLI
Natural Language Inference
GNB
Gaussian Naive Bayes
DT
Decision Tree
SVM
Support-Vector Machine
RF
Random Forest
XGBoost
eXtreme Gradient Boosting
MLIR
Multilingual Information Retrieval
IR
Information Retrieval
NDCG
Normalized Discounted Cumulative Gain
LD
Leading Decision
FSCD
Federal Supreme Court Decisons
SFCS
Swiss Federal Supreme Court
LLM
Large Language Model
LM
Language Model
CVG
Court View Generation
JP
Judgment Prediction
LAP
Law Area Prediction
SLAP
Sub Law Area Prediction
CP
Criticality Prediction
LDS
Leading Decision Summarization
CE
Citation Extraction
SBERT
Sentence-Bert
Doc2Doc
Document-to-Document
mUSE
Multilingual Universal Sentence Encoder
1 Introduction
Predicting the impact of legal cases is a critical task in the legal domain, as it aids professionals in the judicial system in navigating large volumes of case law. Despite its significance, the task of predicting the case criticality remains relatively under-explored. Existing approaches to evaluating the importance of legal cases are primarily manual, very resource-intensive and subject to the judgments of individual annotators. This paper introduces a novel dataset, licensed under CC BY 4.0, [^1] https://huggingface.co/datasets/rcds/swiss_criticality_prediction and a more challenging evaluation framework—Criticality Prediction—designed to predict the potential influence of Swiss Federal Court cases on future jurisprudence.
While prior work such as the Importance Prediction task proposed by Chalkidis et al. (2019) for European Court of Human Rights (ECtHR) cases focused on predicting importance on a defined scale using human-assigned labels, our approach employs algorithmically derived labels to evaluate case criticality. Our dataset introduces a two-tier labeling system: the LD-Label, a binary indicator of whether a case is published as a Leading Decision (LD), and the Citation-Label, a more nuanced categorization based on the frequency and recency-weighted citation counts of these decisions across subsequent cases. This automated and distinct formulation of "criticality" not only distinguishes critical cases from non-critical ones but also ranks them by their relative importance over time. The complexity of this dataset challenges even recent large language models (LLMs) such as GPT-3.5 (Brown et al., 2020).
Our contributions are threefold: (1) We propose a novel Criticality Prediction task that provides a more comprehensive and challenging evaluation of case law importance. (2) We release the datasets to the community, providing valuable resources for further research in legal NLP. (3) We evaluate several multilingual models of various sizes, including fine-tuned variants to set baselines.
Figure 1: Overview of the Criticality Prediction Task.
2 Related Work
One of the most common text classification tasks in the legal domain is Legal Judgment Prediction (LJP), which involves predicting the outcome based on its facts. Researchers have leveraged diverse datasets with unique characteristics and annotations to analyze and predict case outcomes across various languages, jurisdictions, and input types (Feng et al., 2022; Aletras et al., 2016; Şulea et al., 2017; Medvedeva et al., 2018; Chalkidis et al., 2019; Niklaus et al., 2021, 2022; Semo et al., 2022). While LJP is focused on the outcome of individual cases, Importance Prediction shifts the focus toward assessing the broader significance of a case.
Chalkidis et al. (2019) introduced the Importance Prediction task using cases from the European Court of Human Rights (ECtHR). In this task, ECtHR provided scores that denote each case’s ‘importance’ to the common law. These scores, ranging from 1 (key case) to 4 (unimportant), were designed to help legal practitioners identify cases that play a crucial role in shaping jurisprudence. The labels reflect the long-term impact of a case on future rulings and the evolution of legal precedent. While the task is invaluable for identifying landmark cases, it relies on legal experts to assign the labels, making the process resource-intensive and potentially subject to subjective interpretations. Additionally, the ECtHR dataset is monolingual in English, whereas our Swiss dataset is multilingual. To our knowledge, no other study addresses a similar task to Criticality Prediction.
3 Task and Dataset
Table 1: Task Configurations. Label names are Critical (C), Non-critical (NC), Critical-1 (C1) to Critical-4 (C4).
Task Name Train Labels Train Validation Labels Validation Test Labels Test
C NC C NC C NC
LD-Facts 74799 2542 72257 - - 12019 580 11439 - - 26239 950 25289 - -
LD-Considerations 87555 2544 85011 - - 13386 580 12806 - - 29486 948 28538 - -
C-1 C-2 C-3 C-4 C-1 C-2 C-3 C-4 C-1 C-2 C-3 C-4
Citation–Facts 2506 782 626 585 513 563 186 152 131 94 725 137 177 224 187
Citation-Considerations 2509 779 624 586 520 563 186 154 131 92 723 137 177 224 185
Figure 2: Facts (Blue) and consideration (Orange) length distribution measured in Words with Spacy Honnibal et al. (2020). Those with more than 6000 words are binned together.
3.1 Criticality
Understanding the legal framework of the Swiss Federal Supreme Court (SFSC) is key to defining criticality in Swiss case law. The SFSC shapes the legal landscape through its rulings, with a subset known as Leading Decisions published separately due to their influence on future decisions. We quantify case criticality using two labels: the binary LD-Label and the more granular Citation-Label.
The LD-Label is binary, categorizing cases as critical or non-critical. SFSC cases are labeled critical if also published as LD, reflecting their recognized importance within the Swiss legal system. We used regular expressions to extract SFSC cases from LD headers if present. SFSC cases not published as LD are labeled non-critical.
We developed the Citation-Label to provide a more granular measure. It counts how often each LD case is cited in SFSC cases, with older cases weighted less to prioritize recency. The score is calculated as: $score=count\times\frac{year-2002+1}{2023-2002+1}$. The count is the citation frequency, and the weighting reduces older cases’ influence. More details on the constants are in Appendix A. This score ranks LD cases, and we categorized them into four criticality levels—critical-1 (least critical) to critical-4—based on the 25th, 50th, and 75th percentiles.
Unlike prior approaches such as  Chalkidis et al. (2019), we explicitly incorporate temporal weighting to account for both the influence of a case and how its criticality shifts over time. Furthermore, our framework allows for the dynamic recalculation of scores and re-labeling of criticality as case law evolves. This adaptability ensures that our dataset reflects the ongoing changes in the legal system, whereas prior studies would require manual re-annotation as the law develops.
3.2 Criticality Prediction Task and Dataset
The Criticality Prediction (CP) task leverages two primary inputs from the Swiss Federal Supreme Court (SFCS) cases: facts and considerations. Facts describe a factual account of the events of each case and form the basis for the considerations of the court. Considerations reflect the formal legal reasoning, citing laws and other influential rulings, and forming the basis for the final ruling.
We see two distinct applications for these inputs in the Criticality Prediction task, illustrated in Figure 1. In the Case Prioritization task, only the facts are used as input. This produces a score indicating how critical or important a case is. The goal is to help prioritize cases, which could assist in determining which cases should be heard sooner or assigned to more experienced judges.
In the Leading Decisions Selection task, considerations are used for a post-hoc analysis, comparing the ruling to prior case law to assess its potential impact on future jurisprudence. This analysis is exactly what the Supreme Court does at the end of the year to arrive at the selection of leading decisions.
Our approach, by offering facts and considerations as inputs, reflects different stages of legal processing. Both tasks can utilize the LD-Label or Citation-Label, providing varying levels of granularity. This dual approach enhances the dataset’s practical utility, supporting both early-stage prioritization and post-judgment evaluation, thereby addressing multiple aspects of legal workflows.
Our dataset spans from 2002 to 2023, and was partitioned into training (2002-2015), validation (2016-2017), and test (2018-2022) sets, as outlined in Table 1. We allocated a relatively large test set to accommodate longitudinal studies, including the COVID-19 pandemic years. The dataset consists of 138,531 total cases, with 85,167 in German, 45,451 in French, and 7,913 in Italian. See Figure 2 for length distributions of facts and considerations and Appendix Table 4 for general dataset metadata.
4 Methods
4.1 Models
We evaluated the following models, various size variants: MiniLM (Wang et al., 2020), DistilmBERT (Sanh, 2019), mDeBERTa-v3 (He et al., 2021), XLM-R (Base and Large) (Conneau and Lample, 2019), X-MOD (Base) (Pfeiffer et al., 2022), SwissBERT (Vamvas et al., 2023), mT5 (Small and Base) (Xue et al., 2021), BLOOM (560M) (Scao et al., 2022), Legal-Swiss-RoBERTa and Legal Swiss Longformer (Base) (Rasiah et al., 2023), GPT-3.5 (Brown et al., 2020), and LLaMA-2 (Touvron et al., 2023).
We fine-tuned all models per task, using early stopping on the validation dataset. Due to resource constraints, further fine-tuning of GPT-3.5 and LLaMA-2 is reserved for future work, with their current performance serving as baseline results. SwissBERT and Legal-Swiss models were chosen for their Swiss-specific pretraining, while the other models were chosen for their multilingual capabilities, essential for our multilingual dataset.
We evaluated GPT-3.5 and LLaMA-2 0-shot following Chalkidis (2023), using one instruction and example as input. Samples were randomly selected from the validation set to prevent test set leakage for future evaluations especially for a closed model (GPT-3.5). To manage costs, we limited the validation set to 1000 samples. Our experiments focused solely on zero-shot classification due to the long input lengths. We show the prompts used in Appendix Figure 3 and Figure 4. We used the ChatCompletion API for GPT-3.5 (as of June 7, 2023), and ran LLaMA-2 locally with 4-bit quantization.
4.2 Metrics
We adopt the LEXTREME benchmark setup (Niklaus et al., 2023), and use hierarchical aggregation of macro-averaged F1 scores with the harmonic mean to emphasizes lower scores, promoting fairness across languages and input types. Scores are averaged over random seeds, languages (de, fr, it), and input types (facts or considerations), penalizing models with outlier low scores, encouraging consistent performance across all configurations.
5 Results
Table 2: Configuration aggregate scores. The macro-F1 scores from the language-specific subsets of the test set are provided.
Model LD-F LD-C C-F C-C Agg.
Languages  de  /  fr  /  it  de  /  fr  /  it  de  /  fr  /  it  de  /  fr  /  it  de  /  fr  /  it
MiniLM 57.5 / 53.9 / 52.9 68.1 / 65.4 / 64.2 12.1 / 13.1 / 6.8 24.6 / 21.9 / 17.3 25.7 / 25.7 / 16.7
DistilmBERT 56.3 / 55.6 / 56.8 67.8 / 63.9 / 64.7 20.2 / 18.2 / 20.7 22.6 / 21.6 / 22.2 31.7 / 29.7 / 31.6
mDeBERTa-v3 57.6 / 55.1 / 52.7 73.9 / 68.1 / 67.7 25.4 / 22.8 / 16.8 22.1 / 21.6 / 12.6 34.6 / 32.5 / 23.2
XLM-RBase 59.4 / 56.3 / 56.0 70.2 / 65.4 / 62.5 20.0 / 20.6 / 23.5 26.5 / 22.1 / 23.1 33.7 / 31.5 / 33.4
XLM-RLarge 58.4 / 56.8 / 54.1 70.5 / 67.3 / 66.0 22.5 / 19.7 / 36.2 26.7 / 28.2 / 33.0 46.9 / 33.7 / 43.7
X-MODBase 59.0 / 56.2 / 54.8 71.1 / 68.7 / 64.1 19.8 / 17.2 / 24.4 23.2 / 24.2 / 16.4 32.1 / 30.3 / 29.5
SwissBERT(xlm-vocab) 57.6 / 55.9 / 57.3 72.4 / 69.3 / 61.2 23.8 / 20.3 / 39.4 28.5 / 24.0 / 18.7 36.9 / 32.3 / 35.5
mT5Small 54.8 / 51.7 / 50.3 69.2 / 61.9 / 56.4 14.2 / 16.2 / 10.5 15.9 / 18.1 / 20.2 24.1 / 26.2 / 21.9
mT5Base 54.1 / 52.1 / 50.3 66.4 / 61.9 / 56.8 10.6 / 16.3 / 16.9 18.7 / 18.7 / 22.1 22.1 / 26.6 / 28.2
BLOOM-560m 55.1 / 53.2 / 50.9 64.6 / 65.3 / 56.2 12.6 / 16.1 / 7.1 9.5 / 13.6 / 5.1 18.3 / 23.6 / 10.7
Legal-Swiss-RBase 59.3 / 58.4 / 55.5 73.8 / 69.4 / 68.6 24.3 / 20.5 / 10.5 26.2 / 25.3 / 14.0 36.5 / 33.4, 20.1
Legal-Swiss-RLarge 58.3 / 55.7 / 53.6 71.9 / 68.5 / 66.7 23.0 / 21.3 / 38.7 28.5 / 26.0 / 9.0 36.5 / 33.9 / 23.4
Legal-Swiss-LFBase 60.7 / 58.3 / 55.5 74.8 / 70.0 / 67.8 25.3 / 21.5 / 18.4 29.2 / 26.7 / 9.9 38.6 / 34.7 / 21.3
Table 3: Results using Macro F1, with highest values in bold. ’F’ and ’C’ refer to inputs from Facts or Considerations, while ’LD’ and ’C’ denote LD and Citation labels. Models marked with (*) are zero-shot LLMs evaluated on the validation set.
Model LD-F LD-C C-F C-C Agg.
Random Baseline 36.2 36.0 24.1 25.7 29.5
Majority Baseline 49.1 49.2 11.8 11.8 19.0
MiniLM 54.7 65.8 9.8 20.8 21.8
DistilmBERT 56.2 65.4 19.6 22.1 30.9
mDeBERTa-v3 55.1 69.8 21.0 17.5 29.1
XLM-RBase 57.2 65.9 21.3 23.7 32.8
XLM-RLarge 56.4 67.9 24.4 29.1 37.1
X-MODBase 56.6 67.8 20.0 20.6 30.5
SwissBERT(xlm-vocab) 56.9 67.3 25.7 23.0 34.8
mT5Small 52.2 62.1 13.2 17.9 24.0
mT5Base 52.1 61.5 14.0 19.7 25.4
BLOOM560M 53.0 61.7 10.7 8.0 15.8
Legal-Swiss-RoBERTaBase 57.7 70.5 16.2 20.1 28.0
Legal-Swiss-RoBERTaLarge 55.9 68.9 25.8 16.3 30.2
Legal-Swiss-LFBase 58.1 70.8 21.4 17.4 29.5
GPT-3.5* 46.6 44.8 25.7 16.7 28.1
LLaMA-2* 45.2 26.6 7.0 8.5 12.5
We present results in Table 3, with standard deviations in Appendix Table 5 and scores on the validation dataset in Appendix Table 6. The best performance was achieved by XLM-RLarge, with an aggregate (Agg.) score of 37.1. SwissBERT also demonstrated competitive results, with an Agg. score of 34.8. Interestingly, larger models did not always outperform their smaller counterparts. For example, mT5Base and mT5Small both underperformed compared DistilmBERT in all configurations.
The Legal-Swiss models performed well in LD labels, particularly the Legal-Swiss-LFBase, which achieved the highest scores in LD-F and LD-C. However, their weak performance on Citation labels highlights the dataset’s complexity, even for domain-specific models.
LLMs such as GPT-3.5 and LLaMA-2 underperform fine-tuned models, underlining the need for specialized models for these tasks. The difference is largest in the LD labels where the small fine-tuned models always outperformed LLMs.
Table 2 shows more detailed results on the language specific scores. SwissBERT pretrained with a focus on German achieved the highest aggregate score in German, but interestingly with scores in Italian being the highest by far in C-F. Models pretrained on CC100 (Conneau et al., 2020) (MiniLM, mDeBERTa, XLM-R and X-MOD) exhibited mixed results in French and Italian, with all models performing best in German. MiniLM, mDeBERTa, and X-MOD showed underperformance in Italian but stronger results in French. In contrast, XLM-R, particularly the large variant, demonstrated robust performance in Italian. mT5 models performed well in French, and the base variant additionally also performed well on Italian. BLOOM was much better in French than in other languages, not surprising given it did not have German and Italian in the pretraining data.
Overall, there only seems to be a weak trend connecting higher percentage of a given language in the pretraining corpus leading to better downstream results in that language.
6 Conclusions and Future Work
This work introduced a novel Criticality Prediction task to assess the potential influence of Swiss Federal Court cases on future jurisprudence. Our approach utilizes algorithmically derived labels for a more comprehensive and challenging multilingual evaluation of case law importance compared to existing methods. We also released its multilingual dataset to advance research in legal NLP.
We conducted a comprehensive evaluation of our proposed Criticality Prediction task, comparing a diverse range of models, from smaller multilingual models to large language models like GPT-3.5. Our findings demonstrate that fine-tuned models consistently outperform their zero-shot LLMs, achieving superior macro-F1 scores. This underscores the importance of task-specific adaptation for optimizing performance in legal NLP applications.
Future studies could explore application of the Criticality Prediction task in other legal contexts by incorporating sources from different jurisdictions and languages. This would broaden the impact of this research, providing valuable insights into cross-jurisdictional case influences and enhancing the model’s adaptability to various legal systems.
Limitations
It is very difficult to estimate the importance of a case. By relying on proxies such as whether the case was converted to a leading decision (LD-label) and how often this leading decision was cited (Citation-label), we were able to create labels semi-automatically. While we discussed this with lawyers at length and implemented the solution we agreed on finally, this task remains somewhat artificial.
Ethics Statement
While automating case prioritization and identifying leading decisions can greatly benefit legal professionals, there are potential risks associated with deploying such classifiers. One concern is the risk of perpetuating biases present in historical legal trends. For instance, case prioritization decisions should not be influenced by factors such as gender, race, or other protected characteristics. We acknowledge these concerns and will pursue measures to mitigate such biases in future work.
Additionally, there are challenges with reproducibility when using closed models like ChatGPT. Since the internal workings of these models are not fully transparent, results may be difficult to replicate. To promote open science, we have provided comprehensive evaluations of open source multilingual models, aiming to make our findings more accessible and reproducible.
Appendix A Weighting Formula for Citation-Label
The weighting formula used for the Citation-Label is designed to balance the impact of older cases with more recent rulings, ensuring that the ranking reflects both citation frequency and recency. The formula is as follows:
$score=count\times\frac{year-2002+1}{2023-2002+1}$
Where:
  • count refers to the number of times a particular case is cited in Swiss Federal Supreme Court (SFSC) decisions.
  • The year 2002 is the starting year of our dataset.
  • The year 2023 is the end point of the our dataset.
We have +1 adjustment in weighting factor $\frac{year-2002+1}{2023-2002+1}$. This ensures that cases from the year 2002 are still included in the weighted calculation and do not receive a weight of zero.
Appendix B Dataset Example
Input
[Consideraions]: Erwägungen: 1. Angefochten ist der in einem kantonal letztinstanzlichen Scheidungsurteil festgesetzte nacheheliche Unterhalt in einem Fr. 30’000.– übersteigenden Umfang; auf die Beschwerde ist somit einzutreten (Art. 72 Abs. 1, Art. 74 Abs. 1 lit. b, Art. 75 Abs. 1 und Art. 90 BGG). 2. Die Parteien pflegten eine klassische Rollenteilung, bei der die Ehefrau die Kinder grosszog und sich um den Haushalt kümmerte. Infolge der Trennung nahm sie im November 2005 wieder eine Arbeitstätigkeit auf und erzielt mit einem 80%-Pensum Fr. 2’955.– netto pro Monat. Beide kantonalen Instanzen haben ihr jedoch auf der Basis einer Vollzeitstelle ein hypothetisches Einkommen von Fr. 3’690.– angerechnet. Das Obergericht hat zwar festgehalten, der Ehefrau sei eine Ausdehnung der Arbeitstätigkeit kaum möglich, gleichzeitig aber erwogen, es sei nicht ersichtlich, weshalb sie nicht einer Vollzeitbeschäftigung nachgehen könne. Ungeachtet dieses Widerspruches wird das Einkommen von Fr. 3’690.– von der Ehefrau ausdrücklich anerkannt, weshalb den nachfolgenden rechtlichen Ausführungen dieser Betrag zugrunde zu legen ist. […]
Metadata: Decision ID: 65aad3f6-33c2-4de2-91c7-436e8143d6ea
Year: 2007
Language: German
Law Area: Civil
LD Label: Critical
Citation Label: Citation-1
Court: CH_BGer
Chamber: CH_BGer_005,
Canton: CH
Region: Federation
  Target: critical-1
Possible LD label: critical, non-critical, Possible citation label: critical-1, critical-2, critical-3, critical-4
Appendix C General Dataset Metadata
Table 4: Listing of cantons, courts, chambers, law-areas
Metadata Number Examples
Cantons 1 Federation (CH)
Courts 1 Supreme Court
Chambers 13 CH-BGer-011 CH-BGer-004 CH-BGer-008 CH-BGer-002 CH-BGer-005 CH-BGer-001 CH-BGer-006 CH-BGer-009 CH-BGer-015 …
Law-Areas 4 Civil, Criminal, Public, Social
Languages 5 German, French, Italian
Appendix D Dataset Licensing
The original case data is available from the Swiss Federal Supreme Court [^2] https://www.bger.ch/de/index.htm and the Entscheidsuche portal [^3] https://entscheidsuche.ch/ was used to download HTML files for each case.
In compliance with the Swiss Federal Supreme Court’s licensing policy [^4] https://www.bger.ch/files/live/sites/bger/files/pdf/de/urteilsveroeffentlichung_d.pdf, we are releasing the dataset under a CC-BY-4.0 license. The link to the dataset will be made available upon acceptance.
Personally identifying information has already been anonymized by the Swiss Federal Supreme Court in accordance with its anonymization rules [^5] https://www.bger.ch/files/live/sites/bger/files/pdf/Reglemente/Anonymisierungsregeln_2020_def__d.pdf.
Appendix E Zero-shot Prompts
Figure 3: Prompt used for Criticality Prediction (LD Facts/Consideration). The LLM is tasked with predicting whether a Swiss Federal Supreme Court decision is labeled as critical or non-critical based on the provided facts or considerations
Criticality Prediction (CP) LD Facts/Consideration
Given the {facts/considerations} from the following Swiss Federal Supreme Court Decision:
{INPUT FROM THE VALIDATION SET}
Federal Supreme Court Decisions in Switzerland that are published additionally get the label critical, those Federal Supreme Court Decisions that are not published additionally, get the label non-critical. Therefore, there are two labels to choose from:
- critical
- non-critical
The relevant label in this case is:
Figure 4: Prompt used for Criticality Prediction (Citation Facts/Consideration). The LLM is tasked with predicting how likely a Swiss Federal Supreme Court decision is to be cited, using a four-tiered label (critical-1 to critical-4), based on the provided facts or considerations.
Criticality Prediction (CP) Citation Facts/Consideration
Given the {facts/considerations} from the following Swiss Federal Supreme Court Decision:
{INPUT FROM THE VALIDATION SET}
How likely is it that this Swiss Federal Supreme Court Decision gets cited. Choose between one of the following labels (a bigger number in the label means that the court decision is more likely to be cited):
- critical-1
- critical-2
- critical-3
- critical-4
The relevant label in this case is:
Appendix F Hyperparameters and Package Settings
We used a fixed learning rate of 1e-5 without tuning, running each experiment with three random seeds (1-3) and excluding seeds with high evaluation losses. Gradient accumulation was applied when GPU memory was insufficient to maintain a final batch size of 64. Training employed early stopping with a patience of 5 epochs, based on validation loss. To reduce costs, AMP mixed precision was used where it didn’t cause overflows (e.g., mDeBERTa-v3). The max-sequence length was set at 2048 for Facts and 4096 for Considerations.
For the analysis of consideration and fact lengths, we used SpaCy’s en_core_web_sm for tokenization.
Appendix G Resource
The experiments were run on NVIDIA GPUs, including the 24GB RTX3090, 32GB V100, 48GB A6000, and 80GB A100, using approximately 50 GPU days in total.
Appendix H Additional Results
Table 5: Configuration aggregate scores with standard deviations on the test set. The macro-F1 scores are provided.
Model LD-F LD-C C-F C-C Agg.
MiniLM 54.7+/-1.9 65.8+/-1.6 9.8+/-2.8 20.8+/-3.0 21.8
DistilmBERT 56.2+/-0.5 65.4+/-1.7 19.6+/-1.1 22.1+/-0.4 30.9
mDeBERTa-v3 55.1+/-2.0 69.8+/-2.8 21.0+/-3.6 17.5+/-4.4 29.1
XLM-RBase 57.2+/-1.5 65.9+/-3.2 21.3+/-1.5 23.7+/-1.9 32.8
XLM-RLarge 56.4+/-1.8 67.9+/-1.9 24.4+/-7.2 29.1+/-2.7 37.1
X-MODBase 56.6+/-1.8 67.8+/-2.9 20.0+/-3.0 20.6+/-3.5 30.5
SwissBERT(xlm-vocab) 56.9+/-0.7 67.3+/-4.7 25.7+/-8.3 23.0+/-4.0 34.8
mT5Small 52.2+/-1.9 62.1+/-5.2 13.2+/-2.4 17.9+/-1.7 24.0
mT5Base 52.1+/-1.6 61.5+/-3.9 14.0+/-2.8 19.7+/-1.6 25.4
BLOOM-560m 53.0+/-1.7 61.7+/-4.1 10.7+/-3.7 8.0+/-3.5 15.8
Legal-Swiss-RBase 57.7+/-1.6 70.5+/-2.3 16.2+/-5.8 20.1+/-5.6 28.0
Legal-Swiss-RLarge 55.9+/-2.2 68.9+/-2.1 25.8+/-7.8 16.3+/-8.7 30.2
Legal-Swiss-LFBase 58.1+/-2.1 70.8+/-2.9 21.4+/-2.9 17.4+/-8.6 29.5
Appendix I Validation Set Result
Table 6: Configuration aggregate scores on the validation set. The macro-F1 scores are provided. The highest values are in bold. It is important to note that the scores presented here are calculated as the harmonic mean over multiple seeds.
Model LD-F LD-C C-F C-C Agg.
MiniLM 59.1 71.0 14.9 36.9 31.9
DistilmBERT 59.6 70.1 26.3 35.8 41.2
mDeBERTa-v3 60.1 73.0 30.4 36.0 44.0
XLM-RBase 60.1 70.5 26.9 38.5 42.6
XLM-RLarge 60.5 71.7 27.2 39.7 43.3
X-MODBase 57.1 71.0 27.0 33.4 40.6
SwissBERT(xlm-vocab) 59.0 72.1 29.4 38.8 44.1
mT5Small 54.8 66.1 26.3 32.5 39.2
mT5Base 55.7 64.4 24.3 29.3 36.8
BLOOM-560m 52.2 64.3 20.1 21.8 30.7
Legal-Swiss-RBase 61.2 73.6 27.7 41.0 44.2
Legal-Swiss-RLarge 61.8 73.5 29.8 32.0 42.3
Legal-Swiss-LFBase 59.4 72.7 32.2 42.5 47.0
Appendix J Use of AI Assistants
We used ChatGPT and Claude to enhance grammatical correctness and style, and utilized Google Colab’s Generate AI feature for some of the dataset analysis.